4. Calculate the concentration of each ions remaining in the solution when 100.0mL of 0.20M Silver Nitrate solution reacts with 100.0mL of 0.15M Calcium Chloride solution. A. 0.050M CI,0.075M Ca ,0.100M NO. B. 0.075M CL,0.050M Ca²+ ,0.100M NO₂. C. 0.100M CI,0.075M Ca² ,0.050M NO


1
Expert's answer
2022-02-09T23:31:02-0500

Concentration of Ca+ = 0.015x1000/200 =0.075M

Concentration of NO3- ions = 0.02x1000/200 =0.1M

Concentration of Cl- ions = 0.01x1000/200 = 0.05M

Therefore, the correct choice is A